# Chain-of-Dictionary Prompting Elicits Translation in Large Language Models

### v5

@@ Abstract
Large language models (LLMs) have shown surprisingly good performance in multilingual neural machine translation (MNMT) even if not being trained explicitly for translation. Yet, they still struggle with translating low-resource languages. As supported by our experiments, a bilingual dictionary between the source and the target language could help. Motivated by the fact that multilingual training effectively improves cross-lingual performance, we show that a chained multilingual dictionary with words expressed in more languages can provide more information to better enhance the LLM translation. To this end, we present a novel framework, CoD, Chain-of-Dictionary Prompting, which augments LLMs with prior knowledge with the chains of multilingual dictionaries for a subset of input words to elicit translation abilities for LLMs. Experiments indicate that ChatGPT and InstructGPT still have room for improvement in translating many language pairs. And CoD elicits large gains by up to 13x chrF++ points for MNMT (3.08 to 42.63 for English to Serbian written in Cyrillic script) on FLORES-200 full devtest set. We demonstrate the importance of chaining the multilingual dictionaries, as well as the superiority of CoD to few-shot in-context learning for low-resource languages. Using CoD helps ChatGPT to obviously surpass the SOTA translator NLLB 3.3B. [^1] Code and data to be released under publication.
@@ 6 Conclusions
CoD is a novel framework that uses chained multilingual dictionaries when prompting large language models (LLMs) for MNMT. We evaluate ChatGPT, InstructGPT, and BLOOM on the FLORES-200 dataset for MNMT. We found that ChatGPT and InstructGPT still have room for improvement in translating many language pairs. CoD elicits large gains by up to 13x chrF++ points for MNMT (3.08 to 42.63 for English to Serbian written in Cyrillic script) on FLORES-200 full devtest set. We also verified the necessity of the chained multilingual dictionaries, and we found that both of them are quite important to CoD. CoD also outperforms few-shot demonstrations which struggle to retrieve relevant demonstrations for low-resource settings. CoD can even surpass the strong SOTA NLLB translator in translation. These facts validate the usefulness of CoD which uses multilingual dictionaries as an external tool that boosts the translation performance of LLMs.

### v6

@@ Abstract
Large language models (LLMs) have shown surprisingly good performance in multilingual neural machine translation (MNMT) even if not being trained explicitly for translation. Yet, they still struggle with translating low-resource languages. As supported by our experiments, a bilingual dictionary between the source and the target language could help. Motivated by the fact that multilingual training effectively improves cross-lingual performance, we show that a chained multilingual dictionary with words expressed in more languages can provide more information to better enhance the LLM translation. To this end, we present a novel framework, CoD, Chain-of-Dictionary Prompting, which augments LLMs with prior knowledge with the chains of multilingual dictionaries for a subset of input words to elicit translation abilities for LLMs. Experiments indicate that ChatGPT and InstructGPT still have room for improvement in translating many language pairs. And CoD elicits large gains by up to 13x chrF++ points for MNMT (3.08 to 42.63 for English to Serbian written in Cyrillic script) on FLORES-200 full devtest set. We demonstrate the importance of chaining the multilingual dictionaries, as well as the superiority of CoD to few-shot in-context learning for low-resource languages. Using CoD helps ChatGPT to obviously surpass the SOTA translator NLLB 3.3B. [^1] Code and resources available at https://github.com/HongyuanLuke/Chain-of-Dictionary.
@@ 4.5 Ablation Study
Figure 3: A case study on translating from English into Kikongo with Latin script using GPT-4 throughout the cases. We evaluate the results on BLEU and chrF++. We highlight in green the words translated wrong by baselines but translated correctly by CoD, even if the words are not presented in the multilingual dictionary chains.
@@
4.6 Case Study
Figure 3 presents a case study demonstrating the powerfulness of CoD. The baseline output from GPT4 is almost lost about which topics are discussed in the sentence. Using a bilingual dictionary is useful, but the bilingual baseline is still lost about the detailed semantics. In comparison, CoD successfully provides a high-quality translation, scoring the best in BLEU and chrF++. We also highlight in green where the translation is successfully elicited by CoD, even if the words are not provided in the multilingual dictionary. We hypothesise that CoD provides richer context to the LLMs to translate relevant words in the source sentences, even if they are not directly presented by CoD. Figure 4 and Figure 5 demonstrate cases that show a similar phenomenon, and they are available in the Appendix, at the end of this paper.
@@ 6 Conclusions
CoD is a novel framework that uses chained multilingual dictionaries when prompting large language models (LLMs) for MNMT. We evaluate ChatGPT, InstructGPT, and BLOOM on the FLORES-200 dataset for MNMT. We found that ChatGPT and InstructGPT still have room for improvement in translating many language pairs. CoD elicits large gains by up to 13x chrF++ points for MNMT (3.08 to 42.63 for English to Serbian written in Cyrillic script) on FLORES-200 full devtest set. We also verified the necessity of the chained multilingual dictionaries, and we found that both of them are quite important to CoD. CoD also outperforms few-shot demonstrations which struggle to retrieve relevant demonstrations for low-resource settings. CoD can even surpass the strong SOTA NLLB translator in translation. Extensive case studies demonstrate that CoD elicits translation even if the words are not directly presented by CoD.
@@ Appendix C InstructGPT
Figure 4: A case study on translating from English into Central Kurdish with Latin script using GPT-4 throughout the cases. We evaluate the results on BLEU and chrF++. We highlight in green the words translated wrong by baselines but translated correctly by CoD, even if the words are not presented in the multilingual dictionary chains.
Figure 5: A case study on translating from English into Central Kurdish with Latin script using GPT-3.5 throughout the cases. We evaluate the results on BLEU and chrF++. We highlight in green the words translated wrong by baselines but translated correctly by CoD, even if the words are not presented in the multilingual dictionary chains.